The temperature cofficient of copper is 0.004°c^-1. Find the

resistance of a 5 m long copper wire of diameter 0.2mm at

100°c it the resistivity of copper at 0ºc is 1.7 x 10^-8 ohm meter.​


1
Expert's answer
2021-06-01T10:48:37-0400

The resistivity at a certain temperature will be


"\rho_T=\rho(1+\alpha(T-T_0))."

The resistance at 100°C:


"R_T=\frac{\rho_TL}{A}=\frac{4\rho_TL}{\pi d^2}=\frac{4L\rho(1+\alpha(T-T_0))}{\pi d^2},\\\space\\
R_T=3.79\Omega."